Solve the equation using the zero-product property.
(2x+2)(5x-5)=0

Found 2 solutions by mananth, rapaljer:
Answer by mananth(16946)

(2x+2)(5x-5)=0
either of the terms in the bracket can be 0 for the equation to be justified.
either (2x+2)=0
Or
(5x-5)=0
if 2x+2 =0
2x=-2
x=-1
..
..
If 5x-5 =0
5x= 5
x=1
if x=1 OR x=-1 then the left side value is 0


Answer by rapaljer(4671)

When a product of two quantities equals 0, then one of those quantities must equal 0. In this case, there are two solutions.

(2x+2)(5x-5)=0

First solution:
2x+2=0
2x=-2
x=-1

Second solution:
5x-5=0
5x=5
x=1

Dr. Robert J. Rapalje, Retired